### v1

@@ 3.3 Key inequalities for analysis of gradient methods
By (strong) convexity of $H(\cdot)$ and convexity of $\mathcal{U}$, we have
@@ 4 Stochastic Subgradient Method
Let A.1, A.2 hold, and $\rho>\ell$, $\eta\leq 1/\rho$ Then for all $t\geq 0$, we have $\mathbb{E}\left[\left\|x^{t+1}-\hat{x}^{t}\right\|^{2}\mid x^{t}\right]\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+2G_{F}^{2}\eta^{2}.$
@@
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\left(\frac{3\alpha^{2}}{2\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right){\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right]+4\ell\eta^{2}G_{F}^{2}.$
@@
$\qquad+\left(1+\frac{1}{s}\right)\frac{\rho}{2}{\mathbb{E}}\left[\left\|\hat{x}^{t}-z\right\|^{2}\right]+\left(1+s\right)\rho\eta^{2}G_{F}^{2},$
@@
$\mathbb{E}\left[\Phi_{1/\rho}\left(x^{t+1}\right)\right]$$\leq$ $\mathbb{E}\left[F\left(z\right)+\left(1-\frac{\eta\rho}{2}\right)\frac{\rho}{2}\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]+\frac{3}{2\eta}{\mathbb{E}}\left[\left\|\hat{x}^{t}-z\right\|^{2}\right]+2\rho\eta^{2}G_{F}^{2}.$
@@
${\mathbb{E}}\left[\Phi_{1/\rho}(x^{t+1})\right]$$\leq$ $(1-\alpha){\mathbb{E}}\left[F(\hat{x}^{t})\right]+\alpha F(x^{*})+\left(1-\frac{\eta\rho}{2}\right)\frac{\rho}{2}{\mathbb{E}}\left[\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]+2\rho\eta^{2}G_{F}^{2}$
@@
$\leq$ $(1-\alpha){\mathbb{E}}\left[\Phi_{1/\rho}(x^{t})\right]+\alpha F(x^{*})+2\rho\eta^{2}G_{F}^{2}$
@@ 4.1 Hidden Convex Setting
Let C.1, C.2, A.1, A.2 hold with $\mu_{H}=0$, and the set $\mathcal{U}$ be bounded by a diameter $D_{\mathcal{U}}$. Fix $\varepsilon>0$, and set the step-size in (3) as $\eta=\frac{1}{2\ell}\cdot\min\left\{1,\frac{\mu_{c}^{2}\varepsilon^{2}}{24D_{\mathcal{U}}^{2}G_{F}^{2}}\right\}.$ Then for $\rho=2l$, we have $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{\ell D_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{\ell D_{\mathcal{U}}^{4}G_{F}^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right)$ iterations.
@@
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\frac{3D_{\mathcal{U}}^{2}\alpha^{2}}{2\mu_{c}^{2}\eta}+4\ell\eta^{2}G_{F}^{2}.$
@@
$\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{3D_{\mathcal{U}}^{2}\alpha}{2\mu_{c}^{2}\eta}+\frac{4\ell\eta^{2}G_{F}^{2}}{\alpha}\leq\varepsilon,$
where the last step holds by setting $\alpha=\min\left\{\eta\ell,\frac{2\varepsilon\mu_{c}^{2}\eta}{3D_{\mathcal{U}}^{2}},\frac{\sqrt{16\ell}\mu_{c}G_{F}\eta^{\nicefrac{{3}}{{2}}}}{\sqrt{3}D_{\mathcal{U}}}\right\}$ after $T=\frac{1}{\alpha}\log\left(\frac{3\Lambda_{0}}{\varepsilon}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\varepsilon}+\frac{\ell D_{\mathcal{U}}^{4}G_{F}^{2}}{\mu_{c}^{4}\varepsilon^{3}}\right).$
@@ 4.2 Hidden Strongly Convex Setting
Let C.1, C.2, A.1, A.2 hold with $\mu_{H}>0$. Then for $\rho=2\ell$ and any $\eta\leq\frac{1}{2\ell}$, $\alpha\leq\min\left\{\eta\ell,\frac{\eta\mu_{c}^{2}\mu_{H}}{2}\right\}$, we have for all $T\geq 0$ $\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{4\ell\eta^{2}G_{F}^{2}}{\alpha}.$ Fix $\varepsilon>0$, and set the step-size in (15) as $\eta=\min\left\{\frac{1}{2\ell},\frac{\mu_{c}^{2}\mu_{H}\varepsilon}{10\ell G_{F}^{2}}\right\}.$ Then $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{\ell}{\mu_{c}^{2}\mu_{H}}+\frac{\ell G_{F}^{2}}{\mu_{c}^{4}\mu_{H}^{2}}\frac{1}{\varepsilon}\right)$ iterations.
@@
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+4\ell\eta^{2}G_{F}^{2}.$
@@
where the first inequality follows by the first-order characterization of strong convexity and the optimality condition, and the last inequality holds by C.2. Recall that $\Lambda_{T}={\mathbb{E}}\left[F(\hat{x}^{T})-F(x^{*})+\frac{\rho}{2}\left\|\hat{x}^{T}-x^{T}\right\|^{2}\right]$ with $\rho=2\ell$. Then
@@ 6 Projected SGD with Momentum
where $(i)$ follows from (19), $(ii)$ holds by Young’s inequality, i.e., $\langle a,b\rangle\leq\frac{\eta}{2}\left\|a\right\|^{2}+\frac{1}{2\eta}\left\|b\right\|^{2}$ with $a=\nabla F(x^{t})-g^{t}$, $b=x^{t+1}-z$, $(iii)$ holds by the smoothness of $F(\cdot)$, i.e., $-\frac{L}{2}\left\|z-x^{t}\right\|^{2}\leq F(x^{t})-F(z)-\langle\nabla F(x^{t}),z-x^{t}\rangle$. We are now ready to utilize the properties of hidden convex functions to bound $F(z)$ and $\left\|z-x^{t}\right\|^{2}$ for some specific choice of $z\in\mathcal{X}$. We select $z:=x_{\alpha}^{t}=c^{-1}((1-{\alpha})c(x^{t})+{\alpha}c(x^{*}))\in\mathcal{X}$, for some $\alpha\in[0,1]$, and $x^{*}\in\mathcal{X}^{*}$. By Proposition 3, we have for $\mu_{H}\geq 0$,
@@ 6.1 Hidden Convex Setting
[…]
@@ 6.2 Hidden Strongly Convex Setting
[…]
@@ Appendix A Technical Lemma
=$\overset{(i)}{=}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-(1-\eta\rho)\eta\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle+\eta^{2}{\mathbb{E}}\left[\left\|g(x^{t},\xi^{t})-\hat{g}^{t}\right\|^{2}\mid x^{t}\right]$
$\overset{(ii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-(1-\eta\rho)\eta\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle+2G_{F}^{2}\eta^{2}$
$\overset{(iii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-(1-\eta\rho)\eta\ell\left\|x^{t}-\hat{x}^{t}\right\|^{2}+2G_{F}^{2}\eta^{2}$
$\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+\eta^{2}G_{F}^{2},$
where in $(i)$ we use unbiasedness of the gradient estimator. In $(ii)$, we use Young’s inequality and A.2, $(iii)$ holds by hypo-monotonicity inequality $\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle\geq-\ell\left\|x^{t}-\hat{x}^{t}\right\|^{2}$. The last inequality holds by the choice of $\rho$ and $\eta$square
@@
Let ϕ()italic-$\phi(\cdot)$ be convex and for some $\eta>0$, $x\in\mathcal{X}$, define $x^{+}:=\operatorname*{arg\,min}_{y\in\mathcal{X}}\left\{\phi(y)+\frac{1}{2\eta}\left\|y-x\right\|^{2}\right\}$, then
@@ Appendix B Further Improvements with Mini-batching
Define $x_{+}^{T}:=\Pi_{\mathcal{X}}(x^{T}-\frac{1}{\rho-L}\nabla F(x^{T}))$, $\rho=4L$. Notice that $F(x_{+}^{T})=\Phi(x_{+}^{T})\leq\Phi_{1/\rho}(x^{T})$, where the inequality follows by [77, Poposition 2.5-(i)] with $\gamma:=(\rho-L)^{-1}$. Therefore, Theorem 5 implies that
@@ Appendix C Historical Remarks
The sub-gradient method, its special case, Projected SGD, and their numerous variants have a long history of development since the first works on stochastic approximation appeared in 1950s [73, 49, 9, 18]. Convex optimization. The case of convex $F(\cdot)$ is particularly well documented [63, 1, 37]. Researchers have studied how to deal with convex constraints, proximal operators, general Bregman divergences [64, 5], and leveraging averaging and momentum schemes [70, 76, 38, 55]. In the convex case, the global convergence of gradient methods in the function value, i.e., find $x\in\mathcal{X}$ with ${\mathbb{E}}\left[F(x)-F(x^{*})\right]\leq\varepsilon$ for any $\varepsilon>0$, is naturally possible and the sample complexity required is ${\mathcal{O}}\left(\varepsilon^{-2}\right)$. [^1] For P-SGD under smooth and bounded variance assumptions, A.1’ and A.2’ in section 2, or for SM under Lipschitz continuity and bounded second moment of stochastic sub-gradients, i.e., A.2.

### v2

@@ 3.3 Key inequalities for analysis of gradient methods
By the (strong) convexity of $H(\cdot)$ and the convexity of $\mathcal{U}$, we have
@@ 4 Stochastic Subgradient Method
Let A.1, A.2 hold, and $\rho=2\ell$, $\eta\leq 1/\rho$ Then for all $t\geq 0$, we have $\mathbb{E}\left[\left\|x^{t+1}-\hat{x}^{t}\right\|^{2}\mid x^{t}\right]\leq(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+4G_{F}^{2}\eta^{2}.$
@@
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\left(\frac{3\alpha^{2}}{2\mu_{c}^{2}\eta}-\frac{(1-\alpha)\alpha\mu_{H}}{2}\right){\mathbb{E}}\left[\left\|c(\hat{x}^{t})-c(x^{*})\right\|^{2}\right]+8\ell\eta^{2}G_{F}^{2}.$
@@
$\qquad+\left(1+\frac{1}{s}\right)\frac{\rho}{2}{\mathbb{E}}\left[\left\|\hat{x}^{t}-z\right\|^{2}\right]+2\left(1+s\right)\rho\eta^{2}G_{F}^{2},$
@@
$\mathbb{E}\left[\Phi_{1/\rho}\left(x^{t+1}\right)\right]$$\leq$ $\mathbb{E}\left[F\left(z\right)+\left(1-\frac{\eta\rho}{2}\right)\frac{\rho}{2}\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]+\frac{3}{2\eta}{\mathbb{E}}\left[\left\|\hat{x}^{t}-z\right\|^{2}\right]+4\rho\eta^{2}G_{F}^{2}.$
@@
${\mathbb{E}}\left[\Phi_{1/\rho}(x^{t+1})\right]$$\leq$ $(1-\alpha){\mathbb{E}}\left[F(\hat{x}^{t})\right]+\alpha F(x^{*})+\left(1-\frac{\eta\rho}{2}\right)\frac{\rho}{2}{\mathbb{E}}\left[\left\|\hat{x}^{t}-x^{t}\right\|^{2}\right]+4\rho\eta^{2}G_{F}^{2}$
@@
$\leq$ $(1-\alpha){\mathbb{E}}\left[\Phi_{1/\rho}(x^{t})\right]+\alpha F(x^{*})+4\rho\eta^{2}G_{F}^{2}$
@@ 4.1 Hidden Convex Setting
Let C.1, C.2, A.1, A.2 hold with $\mu_{H}=0$, and the set $\mathcal{U}$ be bounded by a diameter $D_{\mathcal{U}}$. Fix $\varepsilon>0$, and set the step-size in (3) as $\eta=\frac{1}{2\ell}\cdot\min\left\{1,\frac{\mu_{c}^{2}\varepsilon^{2}}{48D_{\mathcal{U}}^{2}G_{F}^{2}}\right\}.$ Then for $\rho=2\ell$, we have $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{\ell D_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{\ell D_{\mathcal{U}}^{4}G_{F}^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right)$ iterations.
@@
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+\frac{3D_{\mathcal{U}}^{2}\alpha^{2}}{2\mu_{c}^{2}\eta}+8\ell\eta^{2}G_{F}^{2}.$
@@
$\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{3D_{\mathcal{U}}^{2}\alpha}{2\mu_{c}^{2}\eta}+\frac{8\ell\eta^{2}G_{F}^{2}}{\alpha}\leq\varepsilon,$
where the last step holds by setting $\alpha=\min\left\{\eta\ell,\frac{2\varepsilon\mu_{c}^{2}\eta}{9D_{\mathcal{U}}^{2}},\frac{\sqrt{32\ell}\mu_{c}G_{F}\eta^{\nicefrac{{3}}{{2}}}}{\sqrt{3}D_{\mathcal{U}}}\right\}$ after $T=\frac{1}{\alpha}\log\left(\frac{3\Lambda_{0}}{\varepsilon}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\varepsilon}+\frac{\ell D_{\mathcal{U}}^{4}G_{F}^{2}}{\mu_{c}^{4}\varepsilon^{3}}\right).$
@@
The global convergence of iterates of SM in function values follows immediately from Corollary 1. By the Lipschitz continuity of $F$, we have $\mathbb{E}F(x^{T})-F(x^{*})\leq\mathbb{E}F(\hat{x}^{T})-F(x^{*})+G_{F}\mathbb{E}\|\hat{x}^{T}-x^{T}\|\leq\varepsilon+G_{F}\sqrt{\varepsilon\ \ell}$. Taking $\varepsilon$ small, we observe that SM converges globally in function value to arbitrary small accuracy in expectation. In what follows, this result is tightened and improved in several aspects when more structures (hidden strong convexity of $F$ or smoothness of $F$) are available.
@@ 4.2 Hidden Strongly Convex Setting
Let C.1, C.2, A.1, A.2 hold with $\mu_{H}>0$. Then for $\rho=2\ell$ and any $\eta\leq\frac{1}{2\ell}$, $\alpha\leq\min\left\{\eta\ell,\frac{\eta\mu_{c}^{2}\mu_{H}}{2}\right\}$, we have for all $T\geq 0$ $\Lambda_{T}$$\leq$ $(1-\alpha)^{T}\Lambda_{0}+\frac{8\ell\eta^{2}G_{F}^{2}}{\alpha}.$ Fix $\varepsilon>0$, and set the step-size in (21) as $\eta=\min\left\{\frac{1}{2\ell},\frac{\mu_{c}^{2}\mu_{H}\varepsilon}{20\ell G_{F}^{2}}\right\}.$ Then $\Lambda_{T}\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{\ell}{\mu_{c}^{2}\mu_{H}}+\frac{\ell G_{F}^{2}}{\mu_{c}^{4}\mu_{H}^{2}}\frac{1}{\varepsilon}\right)$ iterations.
@@
$\Lambda_{t+1}$$\leq$ $(1-\alpha)\Lambda_{t}+8\ell\eta^{2}G_{F}^{2}.$
@@
where the first inequality follows by the first-order characterization of strong convexity and the optimality condition, and the last inequality holds by C.2.
Recall that $\Lambda_{T}={\mathbb{E}}\left[F(\hat{x}^{T})-F(x^{*})+\frac{\rho}{2}\left\|\hat{x}^{T}-x^{T}\right\|^{2}\right]$ with $\rho=2\ell$. Then
@@ 6 Projected SGD with Momentum
where $(i)$ follows from (25), $(ii)$ holds by Young’s inequality, i.e., $\langle a,b\rangle\leq\frac{\eta}{2}\left\|a\right\|^{2}+\frac{1}{2\eta}\left\|b\right\|^{2}$ with $a=\nabla F(x^{t})-g^{t}$, $b=x^{t+1}-z$, $(iii)$ holds by the smoothness of $F(\cdot)$, i.e., $-\frac{L}{2}\left\|z-x^{t}\right\|^{2}\leq F(x^{t})-F(z)-\langle\nabla F(x^{t}),z-x^{t}\rangle$.
We are now ready to utilize the properties of hidden convex functions to bound $F(z)$ and $\left\|z-x^{t}\right\|^{2}$ for some specific choice of $z\in\mathcal{X}$. We select $z:=x_{\alpha}^{t}=c^{-1}((1-{\alpha})c(x^{t})+{\alpha}c(x^{*}))\in\mathcal{X}$, for some $\alpha\in[0,1]$, and $x^{*}\in\mathcal{X}^{*}$. By Proposition 3, we have for $\mu_{H}\geq 0$,
@@ 6.1 Hidden Convex Setting
Let C.1, C.2, A.1’, A.2’ hold with $\mu_{H}=0$, and the set $\mathcal{U}$ be bounded by a diameter $D_{\mathcal{U}}$. Then for any $\eta\leq\frac{\beta}{4L}$, $\beta\in(0,1]$, and $\alpha\leq\frac{\beta}{2}$, we have for any $T\geq 0$ $\Lambda_{T}^{\text{HB}}\leq(1-\alpha)^{T}\Lambda_{0}^{\text{HB}}+\frac{\alpha D_{\mathcal{U}}^{2}}{\mu_{c}^{2}\eta}+\frac{\beta\eta\sigma^{2}}{\alpha},$ where $\Lambda_{t}^{\text{HB}}$ is given by (23). Fix $\varepsilon>0$, and set the parameters of algorithm (22) as $\eta=\frac{\beta}{4L},\quad\beta=\min\left\{1,\frac{\mu_{c}^{2}}{9D_{\mathcal{U}}^{2}\sigma^{2}}\varepsilon^{2}\right\}.$ Then the scheme (22) returns a point $x^{T}\in\mathcal{X}$ with ${\mathbb{E}}\left[F(x^{T})-F(x^{*})\right]\leq\varepsilon$ when $T=\widetilde{\mathcal{O}}\left(\frac{LD_{\mathcal{U}}^{2}}{\mu_{c}^{2}}\frac{1}{\varepsilon}+\frac{LD_{\mathcal{U}}^{4}\sigma^{2}}{\mu_{c}^{4}}\frac{1}{\varepsilon^{3}}\right).$
@@ 6.2 Hidden Strongly Convex Setting
Let C.1, C.2, A.1’, A.2’ hold with $\mu_{H}>0$. Then for any $\eta\leq\frac{\beta}{4L}$, $\beta\in(0,1]$, and $\alpha\leq\min\left\{\frac{\beta}{2},\frac{\mu_{c}^{2}\mu_{H}\eta}{4}\right\}$, we have for any $T\geq 0$ $\Lambda_{T}^{\text{HB}}\leq(1-\alpha)^{T}\Lambda_{0}^{\text{HB}}+\frac{\beta\eta\sigma^{2}}{\alpha},$ where $\Lambda_{t}^{\text{HB}}$ is given by (23). Fix $\varepsilon>0$, and set the parameters of algorithm (22) as $\eta=\frac{\beta}{4L},\quad\beta=\min\left\{1,\frac{\mu_{c}^{2}\mu_{H}}{8\sigma^{2}}\varepsilon\right\}.$ Then the scheme (22) returns a point $x^{T}\in\mathcal{X}$ with ${\mathbb{E}}\left[F(x^{T})-F(x^{*})\right]\leq\varepsilon$ after $T=\widetilde{\mathcal{O}}\left(\frac{L}{\mu_{c}^{2}\mu_{H}}+\frac{L\sigma^{2}}{\mu_{c}^{4}\mu_{H}^{2}}\frac{1}{\varepsilon}\right).$
@@ Appendix A Technical Lemma
=$\overset{(i)}{=}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-2(1-\eta\rho)\eta\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle+\eta^{2}{\mathbb{E}}\left[\left\|g(x^{t},\xi^{t})-\hat{g}^{t}\right\|^{2}\mid x^{t}\right]$
$\overset{(ii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}-2(1-\eta\rho)\eta\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle+4G_{F}^{2}\eta^{2}$
$\overset{(iii)}{\leq}(1-\eta\rho)^{2}\left\|x^{t}-\hat{x}^{t}\right\|^{2}+2(1-\eta\rho)\eta\ell\left\|x^{t}-\hat{x}^{t}\right\|^{2}+4G_{F}^{2}\eta^{2}$
$=(1-\eta\rho)\left\|x^{t}-\hat{x}^{t}\right\|^{2}+4G_{F}^{2}\eta^{2},$
where in $(i)$ we use unbiasedness of the gradient estimator. In $(ii)$, we use Young’s inequality and A.2, $(iii)$ holds by hypo-monotonicity inequality $\langle g^{t}-\hat{g}^{t},x^{t}-\hat{x}^{t}\rangle\geq-\ell\left\|x^{t}-\hat{x}^{t}\right\|^{2}$. The last equality holds by the choice of $\rho$. $\square$
@@
Let $\phi(\cdot)$ be convex and for any $\eta>0$, $x\in\mathcal{X}$, define
$x^{+}:=\operatorname*{arg\,min}_{y\in\mathcal{X}}\left\{\phi(y)+\frac{1}{2\eta}\left\|y-x\right\|^{2}\right\}.$
Then
@@ Appendix B Further Improvements with Mini-batching
Define $x_{+}^{T}:=\Pi_{\mathcal{X}}(x^{T}-\frac{1}{\rho-L}\nabla F(x^{T}))$, $\rho=4L$. Notice that $F(x_{+}^{T})=\Phi(x_{+}^{T})\leq\Phi_{1/\rho}(x^{T})$, where the inequality follows by [91, Poposition 2.5-(i)] with $\gamma:=(\rho-L)^{-1}$. Therefore, Theorem 5 implies that
@@ Appendix C Historical Remarks
The sub-gradient method, its special case, Projected SGD, and their numerous variants have a long history of development since the first works on stochastic approximation appeared in 1950s [87, 57, 12, 21].
Convex optimization. The case of convex $F(\cdot)$ is particularly well documented [75, 2, 41]. Researchers have studied how to deal with convex constraints, proximal operators, general Bregman divergences [76, 8], and leveraging averaging and momentum schemes [84, 90, 43, 66]. In the convex case, the global convergence of gradient methods in the function value, i.e., find $x\in\mathcal{X}$ with ${\mathbb{E}}\left[F(x)-F(x^{*})\right]\leq\varepsilon$ for any $\varepsilon>0$, is naturally possible and the sample complexity required is ${\mathcal{O}}\left(\varepsilon^{-2}\right)$. [^4] For P-SGD under smooth and bounded variance assumptions, A.1’ and A.2’ in section 2, or for SM under Lipschitz continuity and bounded second moment of stochastic sub-gradients, i.e., A.2.